OpenStudy (anonymous):

Simplify the following expression. square root sign -108

OpenStudy (anonymous):

\[\sqrt{-108?}\]

OpenStudy (anonymous):

Well that can't work unless you want imaginary numbers?

OpenStudy (anonymous):

Can't squareroot a negative number

OpenStudy (anonymous):

Simplify the following expression; express in terms of i. that's what i forgot

OpenStudy (anonymous):

\[6\sqrt{3i}\]

OpenStudy (anonymous):

Does that seem correct or is it maybe\[6i \sqrt{3i}\]

OpenStudy (anonymous):

i have no idea that's why i was asking

OpenStudy (anonymous):

It should be\[6i \sqrt{3}\]

OpenStudy (anonymous):

This is the answer

OpenStudy (anonymous):

okay thanks I really don't like these problems they never really explain it in the book

OpenStudy (anonymous):

I can't understand it myself. How can u have an inaginary number??????